A tennis player returns,with his rackets,a tennis ball of mass 100g coming towards him with a speed of 162 kmh1.If the ball returns with the same speed of and it remains in contact with racket for 0.1s Find the force exerted by him on the ball.Also explain why force is negative.

Grade:9

1 Answers

Khimraj
5 years ago
force = dp/dt
m = 0.1 kg
v = 162*5/18 = 45 m/s
then dp = mv – (-mv) = 2mv = 2*0.1*45 = 9
and dt = 0.1
so F = 9/0.1 = 90 N
Hope it clears.